# What new physics can we extract from inflation using the ACT DR6 and DESI DR2 Observations?

### v2

@@ II Important issues to address
    ×$\quad\quad\times\left|1-\exp\left(-|\Delta N|\frac{\left(n_{s}-1-\frac{r}{8}\right)}{2}\right)\right|,\quad$
@@ III Numerical Analysis
In this section we provide description of our numerical pipeline, designed to compute inflationary observables for a wide array of single–field potentials with a non-minimal coupling to gravity [65, 56, 66]. Our framework accommodates various functional forms of the Jordan-frame potential $V(\phi)$—including monomial, hilltop, plateau, and more exotic shapes—as well as different power-law coupling functions $f(\phi)=(\phi/M_{\rm pl})^{n}$. Emphasis is placed on reproducibility, numerical stability, and flexibility, so that any new potential or coupling choice can be plugged into the same workflow.
III.1 Jordan- to Einstein-Frame Transformation
We begin with the Jordan-frame Lagrangian,
@@ III.1 Jordan- to Einstein-Frame Transformation
where $M_{\rm pl}=2.43\times 10^{18}\,$eV is the reduced Planck mass, and $\xi$ parametrizes the strength of the non-minimal coupling to the Ricci scalar. We typically choose $f(\phi)=(\phi/M_{\rm pl})^{n}$ for integer $n\geq 1$, but the code accepts any smooth $f(\phi)$.
To recast the theory in a form amenable to standard slow-roll analysis, we perform the Weyl rescaling
@@
which renders the gravitational sector purely Einstein–Hilbert but introduces a non-trivial kinetic prefactor for $\phi$. We restore canonical normalization by defining the new field variable
[…]
Numerical evaluation of (25) uses an adaptive Runge–Kutta–Fehlberg (RK45) integrator with relative tolerance $10^{-9}$ and absolute tolerance $10^{-12}$. We discretize $\phi$ on a dense grid of $10^{4}$ points in $[\phi_{\rm end},\phi_{\rm ini}]$, where $\phi_{\rm ini}\sim 30\,M_{\rm pl}$ is chosen to ensure the field begins well within the slow-roll regime. If the integrator detects stiffness or fails to converge, step-size control automatically refines the grid locally.
III.2 Einstein-Frame Potential Construction
@@
We then build $\hat{V}(\varphi)$ by inverting the numerical map $\varphi(\phi)$ via monotonic spline interpolation. Concretely, we fit a cubic spline to the computed $(\phi,\varphi)$ data, evaluate its inverse on a regular $\varphi$ grid, and compute $\hat{V}(\varphi)$ by direct substitution. This two-step procedure maintains an overall accuracy better than $10^{-5}$ in $\hat{V}$.
III.3 Calculation of Slow-Roll Parameters
Working with the canonical field $\varphi$, we compute derivatives of $\hat{V}(\varphi)$ using a five-point central finite-difference stencil. Denoting $d/d\varphi$ by primes, the Hubble slow-roll parameters are
@@ III.3 Calculation of Slow-Roll Parameters
To monitor numerical error, we periodically compare to a three-point stencil; any point exceeding a relative discrepancy of $10^{-3}$ triggers a local refinement of the $\varphi$ grid.
III.4 Termination of Inflation and e-fold Integration
Inflation ends at $\varphi_{\rm end}$ defined by $\epsilon(\varphi_{\rm end})=1$. We bracket this root by scanning for the first grid point satisfying $\epsilon\geq 1$ and then apply Brent’s method to locate $\varphi_{\rm end}$ to within $10^{-4}\,M_{\rm pl}$. The cumulative number of e-folds from $\varphi$ to $\varphi_{\rm end}$ is
@@ III.5 Extraction of Observables
For pivot e-fold values $N_{\star}=50$ and $60$, we invert Eq. (75) to find the corresponding field values $\varphi_{\star}$. At each pivot we then evaluate:
@@
where $g_{*}=106.75$ is the effective count of relativistic degrees of freedom. Uncertainties in $n_{s}$ and $r$ due to numerical tolerances are estimated by propagating a ±$\pm 5\%$ variation in integration tolerances through to the final values.
@@ V.1 The Starobinsky Model
[…]
@@ V.2 Higgs Inflation
[…]
@@ V.3 T-Model Inflation ($m=1$)
[…]
@@ V.5 D-brane Inflation ($p=2$)
[…]
@@
Inflationary Model $f(\phi)=\phi^{k}$ $\xi$ $r$ $n_{s}$ $\alpha_{s}$ $\beta_{s}$ $H_{\mathrm{inf}}$ [$10^{16}$ GeV] $T_{\mathrm{reh}}$ [$10^{16}$ GeV] $\frac{|\Delta\phi|}{M_{\rm Pl}}$
Starobinsky $k=2$ 0.010.010.010.0110101010 0.00299 0.96767 $-5.29\times 10^{-4}$ $2.95\times 10^{-4}$ 0.144 0.320 0.3594
Higgs Inflation $k=2$ 0.010.010.010.0110101010 0.00349 0.96756 $-5.32\times 10^{-4}$ $2.48\times 10^{-4}$ 0.155 0.333 0.3870
T-model ($m=1$) $k=1$ 0.0260.0260.0260.026100100100100 0.00055 0.96669 $-5.55\times 10^{-4}$ $2.13\times 10^{-3}$ 0.0618 0.210 0.1523
Hilltop (Quartic) $k=1$ 0.2220.2220.2220.2229.9839.9839.9839.983 0.00926 0.97352 $-4.51\times 10^{-4}$ $5.32\times 10^{-5}$ 0.253 0.424 0.7046
D-brane Inflation ($p=1$) $k=1$ 1.5021.5021.5021.5021.9971.9971.9971.997 0.00101 0.96309 $-5.91\times 10^{-4}$ $1.29\times 10^{-3}$ 0.0836 0.244 0.1913
@@ VI.1 Starobinsky Model
The running of the spectral index remains slightly negative, with $\alpha_{s}\sim-6\times 10^{-4}$ across the $\xi$ range, while $\beta_{s}$ stabilizes near $+3\times 10^{-5}$. These higher-order observables are consistent with slow-roll expansions where $\alpha_{s}\sim\mathcal{O}(\epsilon^{2})$ and $\beta_{s}\sim\mathcal{O}(\epsilon^{3})$, and their magnitudes suggest negligible small-scale power spectrum distortions. This offers predictive robustness for future CMB missions probing $\mu$-distortions and acoustic reheating.
@@
0.01 0.9628 $4.10\times 10^{-3}$ $-6.00\times 10^{-4}$ 0.390
1 0.9630 $3.90\times 10^{-3}$ $-5.90\times 10^{-4}$ 0.350
10 0.9631 $3.85\times 10^{-3}$ $-5.80\times 10^{-4}$ 0.330
Table 2: Inflationary observables evaluated at $N_{e}=50$ e-folds before the end of inflation for representative values of the non-minimal coupling $\xi$. Running of the running $\beta_{s}\sim(2.8\,\text{\textendash}\,3.0)\times 10^{-5}$ is not shown in the table but included in the numerical analysis.
@@ VII Discriminating Inflationary Models with Forecasted CMB Constraints
Starobinsky $0.961345$ $-7.57\times 10^{-4}$ $0.967668$ $-5.29\times 10^{-4}$
Higgs $0.961187$ $-7.63\times 10^{-4}$ $0.967563$ $-5.32\times 10^{-4}$
D-Brane $0.955997$ $-8.65\times 10^{-4}$ $0.963085$ $-5.91\times 10^{-4}$
Hilltop (Quartic) $0.968125$ $-6.51\times 10^{-4}$ $0.973524$ $-4.51\times 10^{-4}$
$\alpha$-attractor $0.960035$ $-7.99\times 10^{-4}$ $0.966694$ $-5.55\times 10^{-4}$
@@ VIII Scale-Dependence of the Spectral Index: Running and Running of the Running in Nonminimal Inflation
Conversely, Higgs inflation (diamonds) exhibits more complex behavior. While $\alpha_{s}$ remains within $[-6.32,-5.12]\times 10^{-4}$, the second running $\beta_{s}$ demonstrates sharp excursions, such as $\beta_{s}\approx-5.33\times 10^{-3}$ near $\xi\sim 3.36$. These effects are likely driven by radiative corrections from SM gauge bosons and top-quark loops, which alter the effective potential curvature [91]. Such sensitivity introduces theoretical uncertainty and necessitates precision renormalization group analysis to stabilize predictions at high field values.
@@ X.3 CMB B-modes and Tensor Constraints
[…]
@@ X.6 Implications for Inflaton Dynamics
A crucial theoretical consistency relation arises from the Lyth bound [68], linking the tensor amplitude $r$ to the total field displacement $\Delta\phi$ of the inflaton over the last $N$ $e$-folds of inflation:
@@
[…]
@@ XI Model Comparison and Predictions
[…]
@@
These results underscore the power of gravitational wave observables in discriminating among inflationary scenarios that appear degenerate in the $(n_{s},r)$ plane. For example, Starobinsky and $\alpha$-attractors with small $\alpha$ predict identical $(n_{s},r)$ values but diverge significantly in reheating temperature and GW spectral shape in the mid-frequency regime ($f\sim 10^{-2}$$1$ Hz). These frequencies are within reach of next-generation interferometers such as DECIGO and BBO [117], and space-based detectors like LISA [128].
[…]
@@ XII Conclusion
Hilltop Quartic Inflation, characterized by small-field evolution near a local maximum, exhibits moderate tensor signatures ($r\sim 10^{-3}$) and short field excursions $\Delta\phi\sim 0.4\,M_{\rm Pl}$. Though it is marginally consistent with current constraints at $N=60$, it is increasingly disfavored by future forecast ellipses in the $(n_{s},\alpha_{s})$ plane. Remarkably, it predicts the highest amplitude primordial GW background in the intermediate $f\sim 10^{-2}$$1$ Hz regime, with $\Omega_{\rm GW}(f)$ approaching the sensitivity limits of DECIGO, BBO, and MAGIS [122, 117]. This positions Hilltop inflation as a strong candidate for direct GW detection, especially when accounting for preheating-induced amplification effects.

### v3

@@ II Important issues to address
    $\quad\quad\times\left|1-\exp\left(-|\Delta N|\frac{\left(n_{s}-1-\frac{r}{8}\right)}{2}\right)\right|,\quad$
@@ III Numerical Analysis
In this section we provide description of our numerical pipeline, designed to compute inflationary observables for a wide array of single–field potentials with a non‑minimal coupling to gravity [65, 56, 66]. Our framework accommodates various functional forms of the Jordan‑frame potential $V(\phi)$—including monomial, hilltop, plateau, and more exotic shapes—as well as different power‑law coupling functions $f(\phi)=(\phi/M_{\rm pl})^{n}$. Emphasis is placed on reproducibility, numerical stability, and flexibility, so that any new potential or coupling choice can be plugged into the same workflow.
III.1 Jordan‑ to Einstein‑Frame Transformation
We begin with the Jordan‑frame Lagrangian,
@@ III.1 Jordan‑ to Einstein‑Frame Transformation
where $M_{\rm pl}=2.43\times 10^{18}\,$GeV is the reduced Planck mass, and $\xi$ parametrizes the strength of the non‑minimal coupling to the Ricci scalar. We typically choose $f(\phi)=(\phi/M_{\rm pl})^{n}$ for integer $n\geq 1$, but the code accepts any smooth $f(\phi)$.
To recast the theory in a form amenable to standard slow‑roll analysis, we perform the Weyl rescaling
@@
which renders the gravitational sector purely Einstein–Hilbert but introduces a non‑trivial kinetic prefactor for $\phi$. We restore canonical normalization by defining the new field variable
$\varphi(\phi)\;=\;\int_{\phi_{\rm ini}}^{\phi}\sqrt{\frac{1}{\Omega^{2}(\tilde{\phi})}+\frac{3}{2}\Bigl(\frac{d\ln\Omega^{2}}{d\tilde{\phi}}\Bigr)^{\!2}}\;d\tilde{\phi}$. (25)
Numerical evaluation of (25) uses an adaptive Runge–Kutta–Fehlberg (RK45) integrator with relative tolerance $10^{-9}$ and absolute tolerance $10^{-12}$. We discretize $\phi$ on a dense grid of $10^{4}$ points in $\phi_{\rm end},\phi_{\rm ini}]$, where $\phi_{\rm ini}\sim 30\,M_{\rm pl}$ is chosen to ensure the field begins well within the slow‑roll regime. If the integrator detects stiffness or fails to converge, step‐size control automatically refines the grid locally.
III.2 Einstein‑Frame Potential Construction
@@
We then build $\hat{V}(\varphi)$ by inverting the numerical map $\varphi(\phi)$ via monotonic spline interpolation. Concretely, we fit a cubic spline to the computed $(\phi,\varphi)$ data, evaluate its inverse on a regular $\varphi$ grid, and compute $\hat{V}(\varphi)$ by direct substitution. This two‐step procedure maintains an overall accuracy better than $10^{-5}$ in $\hat{V}$.
III.3 Calculation of Slow‑Roll Parameters
Working with the canonical field $\varphi$, we compute derivatives of $\hat{V}(\varphi)$ using a five‑point central finite‑difference stencil. Denoting $d/d\varphi$ by primes, the Hubble slow‑roll parameters are
@@ III.3 Calculation of Slow‑Roll Parameters
To monitor numerical error, we periodically compare to a three‑point stencil; any point exceeding a relative discrepancy of $10^{-3}$ triggers a local refinement of the $\varphi$ grid.
III.4 Termination of Inflation and e‑fold Integration
Inflation ends at $\varphi_{\rm end}$ defined by $\epsilon(\varphi_{\rm end})=1$. We bracket this root by scanning for the first grid point satisfying $\epsilon\geq 1$ and then apply Brent’s method to locate $\varphi_{\rm end}$ to within $10^{-4}\,M_{\rm pl}$. The cumulative number of e‑folds from $\varphi$ to $\varphi_{\rm end}$ is
@@ III.5 Extraction of Observables
For pivot e‑fold values $N_{\star}=50$ and $60$, we invert Eq. (75) to find the corresponding field values $\varphi_{\star}$. At each pivot we then evaluate:
@@
where $g_{*}=106.75$ is the effective count of relativistic degrees of freedom. Uncertainties in $n_{s}$ and $r$ due to numerical tolerances are estimated by propagating a $\pm 5\%$ variation in integration tolerances through to the final values.
@@ V.1 The Starobinsky Model
$N(\phi)=\frac{1}{M_{\rm pl}^{2}}\int_{\phi_{\rm end}}^{\phi}\frac{V}{V^{\prime}}\,d\phi\simeq\frac{3}{4}e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{\rm pl}}}.$ (57)
@@ V.2 Higgs Inflation
$N=\frac{1}{M_{\rm pl}^{2}}\int_{\varphi_{\rm end}}^{\varphi_{\star}}\frac{\hat{V}}{d\hat{V}/d\varphi}d\varphi\simeq\frac{3}{4}e^{\sqrt{\frac{2}{3}}\frac{\varphi_{\star}}{M_{\rm pl}}},$ (75)
@@ V.3 T-Model Inflation ($m=1$)
These predictions are consistent with CMB measurements for $N_{\star}\sim 50$$60$, yielding $n_{s}\sim 0.967$ and $r\sim 0.003$ for $\alpha=1$ [67].
@@ V.5 D-brane Inflation ($p=2$)
$N_{\star}=\frac{1}{M_{\rm pl}^{2}}\int_{\varphi_{\rm end}}^{\varphi_{\star}}\frac{\hat{V}(\varphi)}{d\hat{V}/d\varphi}\,d\varphi.$ (122)
@@
[…]
Starobinsky $k=2$ 0.010.011010 0.00299 0.96767 $5.29\times 10^{-4}$ $2.95\times 10^{-4}$ 0.144 0.320 0.3594
Higgs Inflation $k=2$ 0.010.011010 0.00349 0.96756 $5.32\times 10^{-4}$ $2.48\times 10^{-4}$ 0.155 0.333 0.3870
T-model ($m=1$) $k=1$ 0.0260.026100100 0.00055 0.96669 $5.55\times 10^{-4}$ $2.13\times 10^{-3}$ 0.0618 0.210 0.1523
Hilltop (Quartic) $k=1$ 0.2220.2229.9839.983 0.00926 0.97352 $4.51\times 10^{-4}$ $5.32\times 10^{-5}$ 0.253 0.424 0.7046
D-brane Inflation ($p=1$) $k=1$ 1.5021.5021.9971.997 0.00101 0.96309 $5.91\times 10^{-4}$ $1.29\times 10^{-3}$ 0.0836 0.244 0.1913
@@ VI.1 Starobinsky Model
The running of the spectral index remains slightly negative, with $\alpha_{s}\sim-6\times 10^{-4}$ across the $\xi$ range, while $\beta_{s}$ stabilizes near $3\times 10^{-5}$. These higher-order observables are consistent with slow-roll expansions where $\alpha_{s}\sim\mathcal{O}(\epsilon^{2})$ and $\beta_{s}\sim\mathcal{O}(\epsilon^{3})$, and their magnitudes suggest negligible small-scale power spectrum distortions. This offers predictive robustness for future CMB missions probing $\mu$-distortions and acoustic reheating.
@@
0.01 0.9628 $4.10\times 10^{-3}$ $6.00\times 10^{-4}$ 0.390
1 0.9630 $3.90\times 10^{-3}$ $5.90\times 10^{-4}$ 0.350
10 0.9631 $3.85\times 10^{-3}$ $5.80\times 10^{-4}$ 0.330
Table 2: Inflationary observables evaluated at $N_{e}=50$ e-folds before the end of inflation for representative values of the non-minimal coupling $\xi$. Running of the running $\beta_{s}\sim(2.8\,\text{–}\,3.0)\times 10^{-5}$ is not shown in the table but included in the numerical analysis.
@@ VII Discriminating Inflationary Models with Forecasted CMB Constraints
Starobinsky $0.961345$ $7.57\times 10^{-4}$ $0.967668$ $5.29\times 10^{-4}$
Higgs $0.961187$ $7.63\times 10^{-4}$ $0.967563$ $5.32\times 10^{-4}$
D-Brane $0.955997$ $8.65\times 10^{-4}$ $0.963085$ $5.91\times 10^{-4}$
Hilltop (Quartic) $0.968125$ $6.51\times 10^{-4}$ $0.973524$ $4.51\times 10^{-4}$
$\alpha$-attractor $0.960035$ $7.99\times 10^{-4}$ $0.966694$ $5.55\times 10^{-4}$
@@ VIII Scale-Dependence of the Spectral Index: Running and Running of the Running in Nonminimal Inflation
Conversely, Higgs inflation (diamonds) exhibits more complex behavior. While $\alpha_{s}$ remains within $6.32,-5.12]\times 10^{-4}$, the second running $\beta_{s}$ demonstrates sharp excursions, such as $\beta_{s}\approx-5.33\times 10^{-3}$ near $\xi\sim 3.36$. These effects are likely driven by radiative corrections from SM gauge bosons and top-quark loops, which alter the effective potential curvature [91]. Such sensitivity introduces theoretical uncertainty and necessitates precision renormalization group analysis to stabilize predictions at high field values.
@@ X.3 CMB B-modes and Tensor Constraints
where $\Delta_{\zeta}^{2}(k)$ is the scalar power spectrum amplitude and $T_{\ell}$ is the tensor transfer function, sharply peaked at horizon crossing near recombination. In plateau-like slow-roll potentials such as Starobinsky or Higgs inflation, the predicted tensor-to-scalar ratio is $r\sim(2.5$$4)\times 10^{-3}$, yielding $C_{\ell}^{BB}\sim 10^{-2}\,\mu\mathrm{K}^{2}$ at the recombination bump [58]. These signals are within the projected sensitivity of LiteBIRD [120], CMB-S4 [59], and the Simons Observatory [121], which aim to reach $\sigma(r)\sim 10^{-4}$ levels.
@@ X.6 Implications for Inflaton Dynamics
A crucial theoretical consistency relation arises from the Lyth bound [68], linking the tensor amplitude $r$ to the total field displacement $\Delta\phi$ of the inflaton over the last NN $e$-folds of inflation:
@@
This inequality implies that large $r$ values—and hence detectable $\Omega_{\rm GW}$ signals—require super-Planckian field excursions ($\Delta\phi>M_{\rm Pl}$ for $r\gtrsim 0.01$, N=50N=506060), characteristic of large-field models like chaotic inflation or $\alpha$-attractors. In contrast, small-field models such as hybrid or hilltop inflation involve sub-Planckian excursions, leading to suppressed GW signatures. Nevertheless, such models remain within reach of ultra-sensitive missions like LiteBIRD and CMB-S4.
@@ XI Model Comparison and Predictions
Hilltop Quartic Inflation: This class of small-field models posits inflation near a local maximum, yielding moderately small tensor amplitudes ($r\sim 10^{-3}$) and $\Delta\phi\sim 0.4\,M_{\rm Pl}$ [75]. Although the potential curvature near the hilltop enhances scalar modes, the tensor sector remains suppressed. The inflationary scale is reduced to $H_{\rm inf}\sim 1.5$$1.7\times 10^{13}$ GeV, but reheating can be moderately efficient with $T_{\rm reh}\sim 3.2$$3.5\times 10^{15}$ GeV, depending on the coupling structure to reheating channels.
@@
These results underscore the power of gravitational wave observables in discriminating among inflationary scenarios that appear degenerate in the $(n_{s},r)$ plane. For example, Starobinsky and $\alpha$-attractors with small $\alpha$ predict identical $(n_{s},r)$ values but diverge significantly in reheating temperature and GW spectral shape in the mid-frequency regime ($f\sim 10^{-2}$11 Hz). These frequencies are within reach of next-generation interferometers such as DECIGO and BBO [117], and space-based detectors like LISA [128].
The redshifted tensor spectrum, $\Omega_{\rm GW}(f)h^{2}$, shown in Figure 8, encapsulates these differences. All models converge to a scale-invariant plateau at high frequencies, modulated by the tensor tilt $n_{t}$ and reheating-induced suppression at characteristic scales $f\sim f_{\rm reh}$. In particular, Hilltop and T-models with moderately high $r$ values exhibit elevated $\Omega_{\rm GW}(f)$ in the $f\sim 0.01$11 Hz regime, approaching sensitivity limits of planned missions. D-brane scenarios, while exhibiting suppressed tensor spectra at CMB scales, can produce comparatively large $\Omega_{\rm GW}$ at higher frequencies due to exotic preheating or brane annihilation dynamics.
@@ XII Conclusion
Hilltop Quartic Inflation, characterized by small-field evolution near a local maximum, exhibits moderate tensor signatures ($r\sim 10^{-3}$) and short field excursions $\Delta\phi\sim 0.4\,M_{\rm Pl}$. Though it is marginally consistent with current constraints at $N=60$, it is increasingly disfavored by future forecast ellipses in the $(n_{s},\alpha_{s})$ plane. Remarkably, it predicts the highest amplitude primordial GW background in the intermediate $f\sim 10^{-2}$11 Hz regime, with $\Omega_{\rm GW}(f)$ approaching the sensitivity limits of DECIGO, BBO, and MAGIS [122, 117]. This positions Hilltop inflation as a strong candidate for direct GW detection, especially when accounting for preheating-induced amplification effects.